core: 0 of 1 pt .4.3 Find the critical values χ2L and χ2R for the given confidence level c and sample size n. 3 of 10 (3 comple c 0.8, n 27 (Round to three decimal places as needed)
Answer #1

Given:

c = Confidence Level = 0.8

So,

\alpha = 0.2

n= sample size = 27

\chi_{L} ^{2}= \chi _{\alpha /2,n-1}^{2}= \chi_{0.1,26} ^{2}=35.563

\chi_{R} ^{2}=\chi _{1-\alpha /2,n-1}^{2}=\chi _{0.9,26}^{2}=17.292

The value of 35.563 is got by referring to Chi square Table for \alpha = 0.1 and Degrees of Freedom = 26.

The value of 17.292 is got by referring to Chi square Table for \alpha = 0.9 and Degrees of Freedom =26